OpenStudy (anonymous):

a 55 kg baseball player slides into third base with an initial speed of 4.6 m/s If the coefficient of kinetic friction between the player and the ground is 0.46, what is the player's acceleration?

OpenStudy (anonymous):

So first you should really draw a free body diagram, but this website isn't letting me So we will move onto using Newton's Laws \[F=m*a\] in the vertical direction \[N-m*g=0\] so \[N=m*g\] in the horizontal direction \[-u*N=m*a\] so \[a=-u*N/(m)\] so \[a=-u*(m*g)/m=-u*g\] It seems that a lot of excess information in this problem but it doesn't actually look that bad

OpenStudy (anonymous):

Alright so after plugging all that in should I have gotten something like \[a=-4.508 m/s ^{2}\]

OpenStudy (anonymous):

Yep, that seems reasonable

OpenStudy (anonymous):

Alright thank you so much! Life saver! Now I can do the rest of this work. thanks again!

OpenStudy (anonymous):

Sure thing man, good luck!